Grade 12Physical Chemistry

The density of mercury is 13.6 g/ml. Calculate approximately the diameter of an atom of
mercury assuming that each atom is occupying a cube of edge length equal to the diameter of the mercury atom.

KUMAR
10 Years agoGrade 12

1 Answer

pralay
ApprovedApproved Tutor Answer10 Years ago
Avagadros number =6.023*10^23
At. wt. of mercury(Hg) = 200
In 1 g of Hg, the total number of atom
= 6.023 *1023/200 = 6.023 *1023/2 *102
 3.0115 *1021 = 3.012 * 1021
Density of Mercury (Hg) = 13.6 g/c.c.
mass of 3.012 * 1021 atoms = 1/3.012 *1021
Now volume of 1 atom of mercury (Hg)
= 1/3.012 *1021 *13.6 c.c. = 103 *10/3012 *1021 * 136 c.c.
= 10-17/3012 *136 c.c. 10-17/409632 c.c. = 1000000 *10-23/409632 c.c.
= 2.44 *1023 c.c.
Since each mercury atom occupies a cube of edge length equal to its diameter, therefore,
Diameter of one Hg atom = (2.44 * 10-23)1/3 cm
= (24.4 * 10-24)1/3 cm.
= 2.905 * 10-8cm
=2.91angstom units